Given there is 0.85 probability that any given adult knows Twitter, find the probability of getting four adults who know of Twitter when seven adults are randomly selected.


1
Expert's answer
2021-05-11T08:59:27-0400

"p=0.85 \\

q = 0.15 \\

n = 7 \\

P(X=4) = C^7_4 (0.85)^4(0.15)^{7-4} \\

= \frac{7!}{4!(7-4)!} \times 0.522 \times 0.003375 \\

= 0.0616"

Answer: 0.0616